A baker sells cookies for $0.65 each and small pies for $1.35 each. She has sold three times as many cookies as she has pies at the end of the day she sold $39.60 worth of cookies and small pies. How many small pies did she sell this day?
Answer by TimothyLamb(4379)

x = number of cookies
y = number of pies
---
0.65x + 1.35y = 39.60
x = 3y
---
put the system of linear equations into standard form
---
0.65x + 1.35y = 39.60
x - 3y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of cookies = 36
y = number of pies = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php